OpenStudy (albert0898):

What is the maximum possible area, in square inches, of a rectangle with a perimeter of 20 inches?

OpenStudy (anonymous):

K.... give me a sec...

myininaya (myininaya):

is this a calculus question @Albert0898

myininaya (myininaya):

you are given 2x+2y=20 or dividing both sides by 2 you are given x+y=10 or if you solve for y you are given y=10-x You want to maximize A=xy replace y with (10-x) Find A' and then the critical numbers

OpenStudy (anonymous):

nice. Got it before me. Guess I was to slow.